Question Library
Browse, filter, and study verified past examination questions with step-by-step mathematical reasoning.
Searching 18,000+ Questions...
Sifting topics, KaTeX formulas & verified solutions
Showing 286–300 of 345 verified questions
Sort by:
Mathematics
UTME / JAMB 1992
Statistics and Probability
Solution
\(\begin{array}{c|c} \text{No. of children} & 0 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline \text{No. of families} & 7 & 11 & 6 & 7 & 7 & 5 & 3 \end{array}\)
Find the mode and median respectively of the distribution above
A
2, 1
B
1, 2
C
1, 5
D
5, 2
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
From the table, the mode = 1.
The median = 2.
Source: Myschool Past Questions Archive (Record #31359)
Mathematics
UTME / JAMB 1992
Statistics and Probability
Solution
Two perfect dice are thrown together, Determine the probability of obtaining a total score of 8
A
\(\frac{1}{12}\)
B
\(\frac{5}{36}\)
C
\(\frac{1}{6}\)
D
\(\frac{7}{36}\)
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
\(\begin{array}{c|c} & & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline a & 1 & 1, 1 & 1, 2 & 1, 3 & 1, 4 & 1, 5 & 1, 6\\\hline & 2 & 1,1 & 2, 2 & 2, 3 & 2, 4 & 2, 5 & 2, 6 \\\hline B & 3 & 3, 1 & 3, 2 & 3, 3 & 3, 4 & 3, 5 & 3, 6\\\hline & 4 & 4 , 1 & 4, 2 & 4, 3 & 4, 4 & 4, 5 & 4, 6\\\hline Die & 5 & 5, 1 & 5, 2 & 5, 3 & 5, 4 & 5, 5 & 5, 6 \\\hline & 6 & 6, 1 & 6, 2 & 6, 3 & 6, 4 & 6, 5 & 6, 6\end{array}\)
Probability of obtaining total score of 8 = \(\frac{5}{36}\)
Source: Myschool Past Questions Archive (Record #31362)
Mathematics
UTME / JAMB 1992
Statistics and Probability
Solution
The probability of an event P is \(\frac{3}{4}\) while that of another event Q is \(\frac{1}{6}\). If the probability of both P and Q is \(\frac{1}{2}\). What is the probability of either P or Q.
A
\(\frac{1}{96}\)
B
\(\frac{1}{8}\)
C
\(\frac{5}{6}\)
D
\(\frac{11}{12}\)
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
Prob of P = \(\frac{3}{4}\)
Prob of Q = \(\frac{1}{6}\)
Prob of both P or Q = \(\frac{3}{4} + \frac{1}{6} = \frac{11}{12}\)
Source: Myschool Past Questions Archive (Record #31364)
Mathematics
WAEC / WASSCE 1992
Statistics and Probability
Solution
(a) In a game, a fair die is rolled once and two unbiased coins are tossed at once. What is the probability of obtaining 3 and a tail?
(b) A box contains 10 marbles, 7 of which are black and 3 are red. Two marbles are drawn one after the other without replacement. Find the probability of getting:
(i) a red, then a black marble ; (ii) two black marbles.
medium Difficulty
Verified Correct Answer: Option
Step-by-Step Mathematical Solution
(a) P(obtaining a 3 from a die) = \(\frac{1}{6}\)
P(obtaining a tail in 2 coins tossed at once) = \(\frac{1}{2}\).
\(\therefore\) P(obtaining 3 and a tail = \(\frac{1}{6} \times \frac{1}{2} = \frac{1}{12}\)
(b) No of black marbles = n(B) = 7
No of red marbles = n(R) = 3
Total marble = 10
(i) P(red then black) = P(1st red, then black) or P(1st black, then red)
= \(\frac{3}{10} \times \frac{7}{9} + \frac{7}{10} \times \frac{3}{9}\)
= \(\frac{7}{15}\)
(ii) Prob (two blacks) = \(\frac{7}{10} \times \frac{6}{9} = \frac{7}{15}\)
Source: Myschool Past Questions Archive (Record #54596)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Solution
The mean of 30 observations recorded in an experiment is 5. lf the observed largest value of 34 is deleted, find the mean of the remaining observations
A
5
B
4
C
3.9
D
3.4
E
3.0
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
x = 5, \(\frac{\sum {fx}}{\sum {f}}\) =
5= \(\frac{\sum {fx}}{30}\)
\(\sum {fx}\) = 5 x 30 = 150; 150 - 34 = 116
\(\sum {f}\) = 29
\(\sum {fx}\) = 116
= \(\frac{116}{29}\) = 4
Source: Myschool Past Questions Archive (Record #22930)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Diagram
Solution
Find the mean of the distribution
Attached Figure / Diagram
A
3.5
B
4.0
C
4.4
D
5.0
E
5
medium Difficulty
Verified Correct Answer: Option C
Step-by-Step Mathematical Solution
\(\bar{x}\) = \(\frac{2 + 14 + 15 + 16 + 45 + 42 + 42}{40}\) = \(\frac{176}{40}\) = 4.4
Source: Myschool Past Questions Archive (Record #22931)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Diagram
Find the mode of the distribution
Attached Figure / Diagram
A
40
B
9
C
7
D
5
E
2
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
No expanded working has been attached for this question yet. The verified answer is Option D.
Source: Myschool Past Questions Archive (Record #22932)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Solution
No. of chairs
1
2
3
4
5
6
7
Frequency
2
7
5
4
9
7
6
Find the median of the distribution
A
9
B
6.5
C
5
D
4.5
E
4
medium Difficulty
Verified Correct Answer: Option C
Step-by-Step Mathematical Solution
Median = 1, 1, 2, 2, 2, 2, 2, 2, 2, 3, 3, 3, 3, 3, 4, 4, 4, 4, 5,| 5, 5, |5, 5, 5, 5, 5, 5, 6, 6, 6, 6, 6, 6, 6, 7, 7, 7, 7, 7, 7,
Median = \(\frac{5 + 5}{2}\) = 5
Source: Myschool Past Questions Archive (Record #22933)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Solution
A box contains 2 white and 3 blue identical marbles. If two marbles are picked at random, one after the other without replacement, what is the probability of picking two marbles of different colors?
A
2/3
B
3/5
C
2/5
D
7/20
E
3/10
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
Total number of marbles = 5; 1st pick = 2/5 2nd pick = 3/4; 3rd pick = 3/5; 4th pick = 2/4
∴ Probability of picking two marbles of different colors = (2/5 x 3/4) + (3/5 + 2/4) = 12/20 = 3/5
Source: Myschool Past Questions Archive (Record #22936)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Solution
Mrs. Jones is expecting a baby. The probability that it will be a boy is 1/2 and probability that the baby will have blue eyes is 1/4. What is the probability that she will have a blue-eyed boy?
A
1/8
B
1/4
C
3/8
D
1/2
E
3/4
medium Difficulty
Verified Correct Answer: Option A
Step-by-Step Mathematical Solution
(1/2)(1/4) = 1/8
Source: Myschool Past Questions Archive (Record #23107)
Mathematics
UTME / JAMB 1991
Statistics and Probability
Solution
Fifty boxes each of 50 bolts, were inspected for the number which were defective. The following was the result
No of defective per box
4
5
6
7
8
9
No. of boxes
2
7
17
10
8
6
The mean and the median of the distribution are, respectively
A
6.7, 6
B
7.6, 5
C
5.7, 87
D
34, 6
medium Difficulty
Verified Correct Answer: Option A
Step-by-Step Mathematical Solution
No of defective(x)
No of boxes (f)
\(fx\)
4
2
8
5
7
35
6
17
102
7
10
70
8
8
64
9
6
54
\(\sum\)
50
333
\(Mean = \frac{\sum fx}{\sum f}\)
= \(\frac{333}{50} = 6.66 \approxeq 6.7\)
The median is the average of the 25th and 26th position = 6.
Source: Myschool Past Questions Archive (Record #31426)
Mathematics
UTME / JAMB 1991
Statistics and Probability
Solution
A crate of soft drinks contains 10 bottles of Coca-cola, 8 of Fanta and 6 of sprite. If one bottle is selected at random, what is the probability that it is NOT a Cocacola bottle?
A
\(\frac{5}{12}\)
B
\(\frac{1}{3}\)
C
\(\frac{3}{4}\)
D
\(\frac{7}{12}\)
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
Coca-cola = 10 bottles, Fanta = 8 bottles
Sprite = 6 bottles, Total = 24
P(cola-cola) = \(\frac{10}{24}\)
P(not coca-cola) = 1 - \(\frac{10}{24}\)
\(\frac{24 - 10}{24}\) = \(\frac{14}{24}\)
= \(\frac{7}{12}\)
Source: Myschool Past Questions Archive (Record #31428)
Mathematics
WAEC / WASSCE 1991
Statistics and Probability
Solution
A man has 9 identical balls in a bag. Out of these, 3 are black, 2 are blue and the remaining are red.
(a) If a ball is drawn at random, what is the probability that it is (i) not blue? (ii) not red?
(b) If 2 balls are drawn at random, one after the other, what is the probability that both of them will be (i) black, if there is no replacement? (ii) blue, if there is a replacement?
medium Difficulty
Verified Correct Answer: Option
Step-by-Step Mathematical Solution
Total number of balls = 9, No of blue = 2, No of black = 3
\(\therefore\) No of red = 4.
(a)(i) P(not blue) = \( 1 - P(blue)\)
= \(1 - \frac{2}{9} = \frac{7}{9}\)
(ii) P(not red) = \(1 - P(red)\)
= \(1 - \frac{4}{9} = \frac{5}{9}\)
(b) Picking two balls;
(i) P(black) with no replacement = \(\frac{3}{9} \times \frac{2}{8} = \frac{1}{12}\)
(ii) P(blue) with replacement = \(\frac{2}{9} \times \frac{2}{9} = \frac{4}{81}\)
Source: Myschool Past Questions Archive (Record #54609)
Mathematics
UTME / JAMB 1990
Statistics and Probability
Solution
The mean of ten positive numbers is 16. When another number is added, the mean becomes 18. Find the eleventh number
A
3
B
16
C
18
D
38
medium Difficulty
Verified Correct Answer: Option D
Step-by-Step Mathematical Solution
Mean of 10 numbers = 16
The total sum of numbers = 16 x 10 = 160
Mean of 11 numbers = 18
Total sum of numbers = 11 x 18
= 198
The 11th no. = 198 - 160
= 38
Source: WAEC / JAMB / NECO Classroom Archive (Record #31485)
Mathematics
UTME / JAMB 1990
Statistics and Probability
Solution
Two numbers are removed at random from the numbers 1, 2, 3 and 4. What is the probability that the sum of the numbers removed is even?
A
\(\frac{2}{3}\)
B
\(\frac{1}{3}\)
C
\(\frac{1}{2}\)
D
\(\frac{1}{4}\)
medium Difficulty
Verified Correct Answer: Option B
Step-by-Step Mathematical Solution
Total pairs: C(4,2) = 6.
Favorable (even sum): both odd (1 way: {1,3}) or both even (1 way: {2,4}) → 2 ways.
P = \(\frac{2}{6}\) = \(\frac{1}{3}\).
Source: WAEC / JAMB / NECO Classroom Archive (Record #31489)